Suppose a family spends their income on two things: education (S) and all other goods, (X)....

  1. Suppose a family spends their income on two things: education (S) and all other goods, (X). The family earns $60,000 a year. A unit of education costs $10,000 and a unit of other goods costs $15,000.

    1. a) Write the equation for the family’s budget constraint.

    2. b) Draw the family’s budget constraint. Put education spending on the x-axis and spending on all other goods on the y-axis.

    3. c) The local government opens a public school, which offers $20,000 worth of education services. Draw the new budget constraint after the school opens.

    4. d) Draw an indifference curve on the graph in c) that illustrates optimal consumption at 4 units of education. Label this point “A”.

    5. e) Suppose the government closes the local public school. How would this change the optimal amount of consumption of education?
Answer #1

a).

Budget constraint is given as:

15000 X + 10000 S = 60000

b).

c).

Public school offers 20000 worth of education (2 units) services. This implies consumer can now have 2 extra units of education.

Budget becomes:

15000 X + 10000 (S - 2) = 60000

d).

With income saved from 2 free units of education, more units of other goods can be consumed.

e).

Without public school,

Budget:

15000 X + 10000 S = 60000

In new optimal bundle, less of both goods is consumed.

With removal of income saved from 2 free units of education, less units of other goods can be consumed.

If you are satisfied with the answer, please provide a positive rating. Feel free to comment in case of queries.

Have a nice day ahead!